NOAA Office of Satellite and Product Operations

NOAA 18 HIRS Subsystem Summary

Component Description Status
CHANNEL 01 Spectral Channel #01, 14.95 micro meters Orange
CHANNEL 02 Spectral Channel #02, 14.71 micro meters Orange
CHANNEL 03 Spectral Channel #03, 14.49 micro meters Orange
CHANNEL 04 Spectral Channel #04, 14.22 micro meters Orange
CHANNEL 05 Spectral Channel #05, 13.97 micro meters Orange
CHANNEL 06 Spectral Channel #06, 13.64 micro meters Orange
CHANNEL 07 Spectral Channel #07, 13.35 micro meters Orange
CHANNEL 08 Spectral Channel #08, 11.11 micro meters Orange
CHANNEL 09 Spectral Channel #09, 9.71 micro meters Orange
CHANNEL 10 Spectral Channel #10, 12.47 micro meters Orange
CHANNEL 11 Spectral Channel #11, 7.33 micro meters Green
CHANNEL 12 Spectral Channel #12, 6.52 micro meters Green
CHANNEL 13 Spectral Channel #13, 4.57 micro meters Green
CHANNEL 14 Spectral Channel #14, 4.52 micro meters Green
CHANNEL 15 Spectral Channel #15, 4.47 micro meters Green
CHANNEL 16 Spectral Channel #16, 4.45 micro meters Green
CHANNEL 17 Spectral Channel #17, 4.13 micro meters Green
CHANNEL 18 Spectral Channel #18, 4.00 micro meters Green
CHANNEL 19 Spectral Channel #19, 3.76 micro meters Green
CHANNEL 20 Spectral Channel #20, 0.69 micro meters Green
ELECTRONICS Instrument Electronics Green
FILTER WHEEL Instrument Filter Wheel Motors Red
SCAN MOTOR Instrument Scan Motor Green

Subsystem History:

Date Time Component Description
03/24/2015 0:00 FILTER WHEEL Filter Wheel Motor had a spike in current about 257.2mA to which shortly returned back to its nominal operating value. Radiometrics saw slight change in pattern with the current spike. There were a few spikes JDAY 80 and JDAY 79 but the current value reached was less.
01/8/2014 23:15 FILTER WHEEL Filter Wheel Motor had a spike in current which shortly returned back to its nominal operating value. Radiometrics saw slight change in pattern with the current spike.
02/05/2009 00:00:00   HIRS data degraded & unsuable for the user community. Engineering health & safety is nominal
02/05/2009 00:00:00 CHANNEL 10 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 02 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 03 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 04 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 05 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 06 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 07 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 08 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 09 NOAA-18 HIRS increased IR Channel Noise
02/05/2009 00:00:00 CHANNEL 01 NOAA-18 HIRS increased IR Channel Noise
01/13/2009 00:00:00   NOAA-18 HIRS increased IR Channel Noise
05/10/2007 14:58:00 SCAN MOTOR HIRS to full scan mode after test completion
05/10/2007 13:10:00 SCAN MOTOR Park HIRS at space view for test
09/28/2006 13:05:00   Command "Position Disable" return to nomral scanning. End lens test #2
09/27/2006 13:15:00   Lens stabikization test #2: Commanded Position Disable then Nadir Position Disable then Nadir Position then Position Disable then IWT Position and then FW housing heater off.
09/25/2006 13:35:00   Commanded Scan mirror to IWT position and turn ON Filter wheel housing heater.
09/14/2006 18:35:00   Command Position Disable, return HIRS mirror to normal scan mode. End HIRS lens stabilization test.
09/13/2006 20:35:00   Command Filter Wheel Housing Heater off
09/11/2006 17:25:00   Command HIRS Filter Housing Heater on (for lens stabilization test #1).
09/11/2006 16:00:00   Command the HIRS mirror to IWT position (for lens stabilization test #1).
04/11/2006 06:10:00 CHANNEL 01 HIRS long wave channels again degraded with negative impacts on imagery.
04/07/2006 22:30:00 CHANNEL 01 HIRS long wave channels improved back to nominal values, imagery improved.
03/15/2006 23:11:00   Return to normal Dwell Channel
03/15/2006 23:10:00   Return HIRS to Full Scan.
03/15/2006 17:56:00   Put HIRS in Space View for 3 orbits.
03/15/2006 17:55:00   TIP Dwell on NHIRSMTI (HIRS Scan Mtr Current), CH245
03/14/2006 12:00:00 CHANNEL 01 HIRS payload again downgraded to YELLOW status because of long wave channel noise which has resulted in Products being unable to utilize this payload.
02/05/2006 04:48:00 CHANNEL 01 HIRS long wave channel noise continues to occur intermittently sometimes lasting for several days even in the absence of solar array moves and other potential perturbations to the loose LWIR lens cover. Channel-1 most affected.
01/04/2006 12:00:00 CHANNEL 01 Because of improvements in the long wave channel noise situation and an understanding of its causes (loose lens cover), N18 HIRS has been upgraded to Green status.
12/15/2005 00:00:00 CHANNEL 01 TOAR board investigation concluded on HIRS long wave channel noise. Problem attributed to loose LWIR lens cover that is highly sensitive to minor spacecraft attitude purturbations such as solar array offset changes. Channel-1 reported as nominal although channels 2, 4, 5, 6, 7, 9 and 10 are reported as somewhat out of spec.
10/13/2005 15:41:00   Return to normal Dwell Channel (bus voltage)
10/13/2005 15:40:00   Return HIRS to Full Scan
10/13/2005 14:01:00   Put HIRS in Space View
10/13/2005 14:00:00   TIP Dwell on NHIRFMTI (HIRS FW Mtr Current), CH261
08/15/2005 12:00:00 CHANNEL 01 HIRS payload and channel-1 component downgraded to YELLOW status because of degraded imagery (ref: TOAR 448) first detected on 6 June 05.
07/12/2005 17:44:00 FILTER WHEEL HIRS FW to low power mode
07/12/2005 16:21:00 SCAN MOTOR HIRS to full scan mode
07/12/2005 14:40:00 FILTER WHEEL HIRS FW motor to High Power
07/12/2005 12:57:00   Park HIRS in Space View
07/05/2005 18:55:00 SCAN MOTOR HIRS to Full Scan Mode
07/05/2005 17:22:00 SCAN MOTOR Park HIRS in Space View
06/27/2005 18:40:00 SCAN MOTOR HIRS to full scan mode
06/27/2005 17:00:00 SCAN MOTOR Park HIRS at space view
06/21/2005 15:00:00 SCAN MOTOR Place HIRS back in full scan mode
06/21/2005 13:10:00 SCAN MOTOR Park HIRS at cold (space) view target
06/20/2005 23:25:00 SCAN MOTOR Place HIRS in full scan mode
06/20/2005 21:40:00 SCAN MOTOR Park HIRS at warm target
06/10/2005 12:00:00   HIRS imagery less than nominal on several channels, under investigation.
06/09/2005 15:15:00 SCAN MOTOR HIRS to full scan mode
06/09/2005 13:35:00 SCAN MOTOR HIRS scanner park at Space View
06/08/2005 22:01:00 SCAN MOTOR HIRS to full scan mode.
06/08/2005 20:21:00 SCAN MOTOR HIRS scanner park at warm target
06/05/2005 15:55:00 CHANNEL 20 On orbit, operational. End HIRS OV period. Subsytem configured for nominal operations.
06/05/2005 15:54:00 CHANNEL 19 On orbit, operational
06/05/2005 15:53:55 CHANNEL 18 On orbit, operational
06/05/2005 15:53:45 CHANNEL 17 On orbit, operational
06/05/2005 15:53:35 CHANNEL 16 On orbit, operational
06/05/2005 15:53:25 CHANNEL 15 On orbit, operational
06/05/2005 15:53:15 CHANNEL 14 On orbit, operational
06/05/2005 15:53:15 CHANNEL 13 On orbit, operational
06/05/2005 15:53:10 CHANNEL 12 On orbit, operational
06/05/2005 15:53:05 CHANNEL 11 On orbit, operational
06/05/2005 15:53:00 CHANNEL 10 On orbit, operational
06/05/2005 15:52:55 CHANNEL 09 On orbit, operational
06/05/2005 15:52:50 CHANNEL 07 On orbit, operational
06/05/2005 15:52:50 CHANNEL 08 On orbit, operational
06/05/2005 15:52:45 CHANNEL 06 On orbit, operational
06/05/2005 15:52:35 CHANNEL 05 On orbit, operational
06/05/2005 15:52:30 CHANNEL 04 On orbit, operational
06/05/2005 15:52:25 CHANNEL 03 On orbit, operational
06/05/2005 15:52:20 CHANNEL 02 On orbit, operational
06/05/2005 15:52:10 CHANNEL 01 On orbit, operational
06/05/2005 15:52:00 ELECTRONICS HIRS electronics turnon, filter motor to normal, calibration on.
06/04/2005 19:12:00   HIRS outgasssing complete. Enabled patch control heater and opened cooler cover door.
05/20/2005 10:22:00 FILTER WHEEL Spacecraft launched with HIRS filter motor powered up. HIRS detectors not yet activated, no imagery available until outgassing completes (NET 5 June)
05/20/2005 10:22:00 SCAN MOTOR Spacecraft launched with HIRS scan motor powered up. HIRS detectors not yet activated, no imagery available until outgassing completes (NET 5 June)